# Representation Learning for Distributional Perturbation Extrapolation

### v1

@@ 5.3 Population-Level Learning Objective
where ^$\widehat{\mathbb{P}}_{e\to h}$ depends on $(\mathbb{P}_{e},\mathbf{l}_{e},\mathbf{l}_{h})$ and the model parameters via (5.5); and $d$ is a measure of dissimilarity between distributions. Here, we use the negative expected energy score for $d$, i.e.,
@@
$d\left(\widehat{\mathbb{P}}_{e\to h},\mathbb{P}_{h}\right)=-\mathbb{E}_{\mathbf{X}_{h}\sim\mathbb{P}_{h}}\left[\operatorname{ES}_{\beta}\left(\widehat{\mathbb{P}}_{e\to h},\mathbf{X}_{h}\right)\right],$ (5.8)
where $\operatorname{ES}_{\beta}$ denotes the energy-score (Gneiting & Raftery, 2007), defined as
[…]
@@
For $\beta\in(0,2)$, it is a strictly proper scoring rule, meaning that the expected energy score $\mathbb{E}_{\mathbf{X}}[\operatorname{ES}_{\beta}(\mathbb{P},\mathbf{X})]$ is maximised if and only if $\mathbf{X}\sim\mathbb{P}$, see § B.1 for details on probabilistic forecasting and scoring rules. Combined with its computational simplicity, this property makes the negative expected energy score a popular loss function for distributional regression (Shen & Meinshausen, 2024b, a). It also directly implies the following result.
@@
The objective in (5.7) is minimised if and only if $\mathbb{P}_{h}=\widehat{\mathbb{P}}_{e\to h}$ for all $e,h\in[M]_{0}$.
The sum in (5.7) also includes the case $h=e$. In this case, enforcing $\mathbb{P}_{e}=\widehat{\mathbb{P}}_{e\to e}$ can be viewed as a marginal reconstruction objective and thus as a weaker constraint than that enforced by $\mathcal{L}_{\text{cond-rec}}$ below.
Conditional Reconstruction Loss.
The conditional reconstruction loss $\mathcal{L}_{\text{cond-rec}}$ is given by the following sum of domain-specific DPA losses proposed by Shen & Meinshausen (2024a),
$\mathcal{L}_{\text{cond-rec}}\left(\widehat{\mathbf{g}},\widehat{\mathbf{f}};\left\{\mathbb{P}_{e}\right\}_{e\in[M]_{0}}\right):=\sum_{e\in[M]_{0}}-\mathbb{E}_{\mathbf{X}_{e}\sim\mathbb{P}_{e}}\left[\operatorname{ES}_{\beta}\left(\widehat{\mathbb{P}}_{\mathbf{X}|\widehat{\mathbf{g}}(\mathbf{X})=\widehat{\mathbf{g}}(\mathbf{X}_{e})},\mathbf{X}_{e}\right)\right],$ (5.10)
where each summand on the RHS of (5.10) involves a conditional energy score (i.e., one whose first argument also depends on $\mathbf{X}_{e}$), see § B.3 for further details and discussion.
5.4 Training
In practice, we do not have access to the true distributions $\mathbb{P}_{e}$, but only to empirical distributions based on the available data in (2.1). Therefore, we approximate expectations w.r.t. $\mathbb{P}_{e}$ and $\mathbb{P}_{h}$ in (5.8), (5.9), and (5.10) with Monte Carlo estimates based on mini-batches of (2.1), and use a single draw of ${\bm{\varepsilon}}$ to approximate expectations w.r.t. $\mathbb{Q}_{\bm{\varepsilon}}$. Moreover, for a fixed encoder, the optimal perturbation matrix is available in closed form and given by the ordinary least squares solution to regressing the domain-specific means $\widehat{\bm{\mu}}^{\mathbf{z}}_{e}:=\frac{1}{N_{e}}\sum_{i=1}^{N_{e}}\widehat{\mathbf{z}}^{\mathrm{pert}}_{e,i}$ on the corresponding perturbation labels $\mathbf{l}_{e}$. We, therefore, write $\widehat{\mathbf{W}}$ as a function of $\widehat{\mathbf{g}}$ and optimise (5.6) with respect to the parameters of $\widehat{\mathbf{g}}$ and $\widehat{\mathbf{f}}$ using stochastic gradient descent (Kingma & Ba, 2015). To avoid encouraging the encoder to capture latent dimensions of maximal variation as in the DPA (Shen & Meinshausen, 2024a) rather than focus on the perturbation-relevant latents $\mathbf{Z}$, we do not update the encoder parameters using (the gradient of) the reconstruction loss.
5.5 Prediction
@@
To sample from the predicted distribution for a new perturbation label $\mathbf{l}_{\mathrm{test}}$, we use our model to compute the synthetic perturbed test latents $\widehat{\mathbf{z}}^{\mathrm{pert}}_{e\to\mathrm{test},i}$ for all $e\in[M]_{0}$ and all $i\in[N_{e}]$ via (5.2), and then sample the corresponding synthetic test observations $\widehat{\mathbf{X}}_{e\to\mathrm{test},i}$ according to (5.3). In other words, our estimate of $\mathbb{P}_{\mathrm{test}}$ is the empirical version of the pooled domain-specific marginal perturbation distributions $\mathbb{P}_{e\to\mathrm{test}}$ induced by our model for $\mathbf{l}_{\mathrm{test}}$, i.e.,
^$\widehat{\mathbb{P}}_{\mathrm{test}}=\frac{1}{M+1}\sum_{e\in[M]_{0}}\widehat{\mathbb{P}}_{e\to\mathrm{test}}.$ (5.11)
6 Relation to Prior Work
Perturbation Modelling.
Our model from § 3 is closely inspired by prior work on modelling combinations of genetic or chemical perturbations, particularly the compositional perturbation autoencoder (CPA; Lotfollahi et al., 2019). CPA additionally includes dose-response curves that allow for a nonlinear element-wise rescaling of the perturbation labels $\mathbf{l}$ and employs an adversarial autoencoder for estimation, see § B.4 for further details. Other prior works such as SVAE+ (Lopez et al., 2023) and SAMS-VAE (Bereket & Karaletsos, 2023) instead fit a generative model to the data by maximising the evidence lower bound (ELBO; a lower bound to the marginal likelihood) using variational inference, typically within the variational autoencoder (VAE) framework (Kingma & Welling, 2014; Rezende et al., 2014). Since computing the ELBO requires evaluating the likelihood $p(\mathbf{x}|\mathbf{z})$, this approach generally comes with parametric assumptions. In contrast, our PDAE approach is based on classical (non-variational) autoencoders and is trained using the energy score, which only requires being able to sample from the stochastic decoder. Both SVAE+ and SAMS-VAE additionally impose a sparsity assumption on the effect of perturbations on the latents. SAMS-VAE adopts the same compositional perturbation model as CPA (and the present work) wherein the effect of perturbations can be decomposed into a sum of effects of elementary perturbations, as captured by the linear form $\mathbf{W}\mathbf{l}$. In contrast, SVAE+ extends work on identifiable VAEs with sparsely changing mechanisms (Lachapelle et al., 2022) and models each perturbation condition separately. Importantly, unlike the present work, none of these prior works on perturbation modelling provide theoretical identifiability or extrapolation guarantees.
Relation to Causal Models.
In the field of causal inference, experimental data resulting from perturbations is modelled via interventions in an underlying causal model. For example, in the structural causal model (SCM; Pearl, 2009) framework, interventions modify a set of assignments, which determine each variable from its direct causes and unexplained noise. In general, our model for the effect of perturbations in latent space (§ 3) differs from how interventions are treated in SCMs. However, as detailed in Appx. C, if we restrict our attention to linear SCMs, then the class of shift interventions can be viewed as a special case of our model with $K={d_{Z}}$ elementary perturbations and a particular choice of perturbation matrix $\mathbf{W}$. In this sense, our approach may be interpreted as causal representation learning (Schölkopf et al., 2021) with a linear latent causal model (Squires et al., 2023; Buchholz et al., 2024) and generalized shift interventions (Zhang et al., 2024). In fact, Zhang et al. (2024) also consider the task of extrapolating to unseen combinations of genetic perturbations from a causal representation learning perspective by modelling elementary perturbations as soft interventions (shifts in a nonlinear causal model) on individual latent variables.
Identifiable Representation Learning for Extrapolation.
Saengkyongam et al. (2024, §4) consider an identifiability problem similar to the one studied in § 4 of the present work and also leverage affine identifiability to derive out-of-support extrapolation guarantees. However, their model (and proof strategy) differs from ours, for example, in that they assume the training support of the perturbation labels to contain a non-empty open subset of $\mathbb{R}^{K}$.
7 Experiments
We present preliminary empirical evidence that our approach can outperform existing methods and baselines at distributional perturbation extrapolation. As we focus on a simple, controlled setting with synthetic data, our current results should be viewed as proof of concept, rather than a comprehensive empirical study.
Methods.
We compare our approach with CPA (Lotfollahi et al., 2023) and the following baselines:
  • Pool All:
    pool all training observations;
  • Pseudobulking:
    pool only training data arising from individual perturbations involved in the combination to be predicted (e.g., pool data from $\mathbf{l}_{1}$ and $\mathbf{l}_{2}$ below for predicting $\mathbf{l}_{\mathrm{test}}=(1,1,0)^{\top}$);
  • Linear Regression:
    linearly regress the environment-specific means of observations $\bm{\mu}^{\mathbf{x}}_{e}\in\mathbb{R}^{d_{X}}$ on $\mathbf{l}_{e}$ and use the resulting model to predict the test mean $\bm{\mu}_{\mathrm{test}}^{\mathbf{x}}$ from $\mathbf{l}_{\mathrm{test}}$.
The former two were also used by Lotfollahi et al. (2023); we propose the latter as an additional baseline. Since CPA and linear regression only predict a mean, in these cases, we apply the test mean shift to all observations in the reference condition to simulate the resulting test distribution and use this to also compute distributional similarity metrics.
Metrics.
To assess the distributional fit between predicted and true (empirical) test distributions, we use the energy distance (ED; Székely & Rizzo, 2013), i.e., twice the normalized negative expected (marginal) energy score, and the maximum mean discrepancy (MMD; Gretton et al., 2012) with Gaussian kernel and bandwidth chosen by the median heuristic, see § B.2 for more on measures of distributional similarity. Since some methods are only desgined to predict the mean, we also report the L2 norm of the difference between the predicted and true mean, $\left\lVert\bm{\mu}^{\mathbf{x}}-\widehat{\bm{\mu}}^{\mathbf{x}}\right\rVert$.
@@ Methods.
Table 1: Results on Synthetic 2D Data. We report average performances over 6 ID and OOD test cases each; shown are the mean $\pm$ one standard deviation over 5 random seeds. Best results highlighted in bold. Theoretical guarantees only apply to ID test cases, i.e., methods cannot be expected to perform well on OOD test cases.
@@
Pool All 0.852 $\pm$ .004 1.124 $\pm$ .006 0.385 $\pm$ .001 2.252 $\pm$ .004 3.222 $\pm$ .008 1.488 $\pm$ .002
Pseudobulking 0.723 $\pm$ .003 1.588 $\pm$ .010 0.441 $\pm$ .001 1.619 $\pm$ .005 3.229 $\pm$ .009 1.201 $\pm$ .003
Linear Regression 0.138 $\pm$ .002 0.613 $\pm$ .010 0.215 $\pm$ .002 0.876 $\pm$ .008 2.865 $\pm$ .026 0.770 $\pm$ .005
CPA 0.339 $\pm$ .126 1.411 $\pm$ .363 0.372 $\pm$ .066 1.404 $\pm$ .177 3.706 $\pm$ .343 1.109 $\pm$ .090
PDAE (Ours) 0.001 $\pm$ .000 0.005 $\pm$ .001 0.016 $\pm$ .003 0.113 $\pm$ .020 0.491 $\pm$ .091 0.231 $\pm$ .021
Data.
@@
For ease of visualisation, we first consider the case of ${d_{Z}}\!=\!{d_{X}}\!=\!2$-dimensional latents and observations. The base distribution $\mathbb{P}_{\mathbf{Z}}$ is a zero-mean, isotropic Gaussian with standard deviation $\sigma\!=\!0.25$. We consider $K\!=\!3$ elementary perturbations with associated shift vectors $\mathbf{w}_{1}\!=\!(1,0)^{\top}$, $\mathbf{w}_{2}\!=\!(0,1)^{\top}$, and $\mathbf{w}_{3}\!=\!(1,1)^{\top}$ and create $M\!+\!1\!=4$ training domains with perturbation labels $\mathbf{l}_{0}\!=\!(0,0,0)^{\top}$, $\mathbf{l}_{1}\!=\!(1,0,0)^{\top}$, $\mathbf{l}_{2}\!=\!(0,1,0)^{\top}$, and $\mathbf{l}_{3}\!=\!(0,0,1)^{\top}$. In this case, the sufficient diversity condition (iii) used to establish affine identifiability in Thm. 4.1 is satisfied. For testing, we distinguish between two scenarios: in-distribution (ID) and out-of-distribution (OOD) test cases, depending on whether the support of the distribution of perturbed latents resulting form $\mathbf{l}_{\mathrm{test}}$ lies (mostly) within or outside the support of perturbed training latents (i.e., the inputs “seen” by the decoder during training), respectively. We consider six ID and OOD test cases each, see Appx. D for a list of the exact test labels $\mathbf{l}_{\mathrm{test}}^{\text{ID}}$ and $\mathbf{l}_{\mathrm{test}}^{\text{OOD}}$. To generate observations, we use the complex exponential $\mathbf{x}\!=\!\mathbf{f}(\mathbf{z})\!=\!\mathrm{e}^{z_{1}}(\cos z_{2},\sin z_{2})$ as a deterministic nonlinear mixing function. The resulting training datasets, along with an example of an ID and an OOD test case are in shown in Fig. 3 (left); see Fig. 4 in Appx. D for a complete visualization including all twelve test cases.
Experimental Details.
@@
We generate $N_{e}=2^{14}$ observations for each domain. Both CPA and PDAE use $4$-hidden-layer MLPs with $64$ hidden units as encoders and decoders and are trained for $2000$ epochs using a batch size of $2^{12}$. For PDAE, we set $\beta=1$ in (5.9) and use a learning rate of $0.005$. Since the ground truth mixing function is deterministic, we use a reconstruction loss weight of $\lambda=0$ and no noise dimensions ($\hat{d_{\epsilon}}=0$). For CPA, we select the learning rates of the autoencoder and adversarial classsifier, as well as the latent bottleneck dimension using a grid search based on performance on a held out portion of the training data, as detailed in Appx. D. All other hyperparameters for CPA are set to their default values. We report results for average performance across the six ID and OOD test cases over five repetitions of randomly generating data and initializing, training, and evaluating models.
@@
Figure 3: Qualitative Comparison of PDAE and CPA on Synthetic 2D Data. Rows correspond to latent space (top) and observation space (bottom). Columns show the ground truth data (left), PDAE predictions (center), and CPA predictions (right). Training domains are shown in grey, an in-distribution (ID) test case with $\mathbf{l}^{\textsc{id}}_{\mathrm{test}}\!=\!(1,1,0)^{\top}$ (which overlaps with one of the training domains) in blue, and an out-of-distribution (OOD) test case with $\mathbf{l}_{\mathrm{test}}^{\textsc{ood}}\!=\!(1,0,1)^{\top}$ in orange. All plots show kernel density estimates of the distributions; crosses (x) indicate the corresponding means. As can be seen, PDAE recovers an affine transformation of the true latents (top, center) leading to accurate distributional predictions for the training and ID test domain (bottom, center). However, the OOD test domain is mapped to a part of the latent space not seen during training (top, center). As a result, the corresponding decoder output does not accurately match the true OOD distribution (bottom left vs center). CPA accurately predicts the means of the training distribution (bottom right, black crosses) but does not recover the true latents up to an affine transformation (top right). Moreover, the predicted distributions for CPA do not match the ground truth particularly well, particularly for the test conditions (bottom left vs right). However, recall that—unlike PDAE—CPA is not trained for distributional reconstruction, see § B.4 for details.
Results.
@@
The quantitative results are summarized in Tab. 1; see also Fig. 3 and its caption for qualitative results. For the ID test setting, PDAE performs best, achieving near-perfect distributional fit. CPA outperforms the pooling and pseudobulking baselines (consistent with results reported by Lotfollahi et al. (2023) on gene perturbation data), but does substantially worse than the linear regression and PDAE at both mean and distributional prediction. For the OOD test setting, all methods perform much worse, with PDAE yielding the least bad performance. This failure for the OOD case is expected. As visualised in Fig. 3, despite learning a good representation and perturbation model (as evident from the strong ID test performance), the decoder did not encounter inputs similar to the perturbed test latents during training and has thus not learnt which part of the observation space to map them to.
Noisy Observations.
@@
In Appx. E, we additionally consider a modified version of the aforementioned experiment with ${d_{X}}=10$ (instead of ${d_{X}}=2$) observed dimensions by concatenating ${d_{\epsilon}}=8$ additional dimensions of mean-zero Gaussian noise and investigate the behaviour for increasing noise variance (i.e., decreasing signal-to-noise ratio). In this case, using a non-zero weight ($\lambda>0$) for the conditional reconstruction loss in (5.10) is helpful and we find PDAE to outperform other methods for low to medium noise variances, see Appx. E for details.
8 Discussion
Single-Cell vs Population-Level Effects.
Since cells are typically destroyed during measurement, each unit is only observed for one perturbation condition. As a result, we do not have access to paired data, which would be required to establish ground-truth single-cell level effects. Instead, we can only observe the effects of perturbations at the population level. Thus, we consider distributional perturbation extrapolation (§ 2) a natural formulation of the objective. Our theoretical guarantees and learning objective are formulated accordingly in distributional terms. This contrasts with some prior works, which, despite a lack of ground truth training data, pursue the (seemingly infeasible) task of making counterfactual predictions at the single-cell level. While our model can, in principle, also make such predictions, we stress that they are not falsifiable from empirical data and that our guarantees do not extend to the single-cell resolution. For this reason, we consider distributional perturbation extrapolation a more meaningful and feasible task formulation.
Is Gaussianity Necessary?
In Thm. 4.1, Gaussianity of $\mathbb{P}_{\mathbf{Z}}$ is assumed to prove identifiability. While this is a sufficient condition, it may not be necessary and can possibly be relaxed. Indeed, our method does not explicitly enforce any particular latent distribution, yet still yields strong performance. However, to encourage normality, we could add a regularizer that penalises the KL between the encoded latent basal state distribution and an isotropic Gaussian prior—akin to the KL loss in VAEs. Moreover, preliminary empirical evidence suggests that training on simulated data generated from a Laplacian basal state distribution also yields identifiable latent spaces and perturbation effects.
Decoder Extrapolation and Uncertainty Quantification.

### v2

@@ 5 Perturbation distribution autoencoder
where $\widehat{\mathbb{P}}_{e\to h}$ depends on $(\mathbb{P}_{e},\mathbf{a}_{e},\mathbf{a}_{h})$ and the model parameters via (5.4); and $d$ measures dissimilarity between distributions. We use the negative expected energy score for $d$,
$d\big(\widehat{\mathbb{P}}_{e\to h},\mathbb{P}_{h}\big)=-\mathbb{E}_{\mathbf{X}_{h}\sim\mathbb{P}_{h}}\big[\operatorname{ES}_{\beta}\big(\widehat{\mathbb{P}}_{e\to h},\mathbf{X}_{h}\big)\big],$ (5.6)
with $\operatorname{ES}_{\beta}$ (Gneiting & Raftery, 2007) defined as
@@
$\operatorname{ES}_{\beta}(\mathbb{P},\mathbf{x})=\frac{1}{2}\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\,\overset{\text{i.i.d.}}{\sim}\,\mathbb{P}}\left\lVert\mathbf{X}-\mathbf{X}^{\prime}\right\rVert^{\beta}-\mathbb{E}_{\mathbf{X}\sim\mathbb{P}}\left\lVert\mathbf{X}-\mathbf{x}\right\rVert^{\beta}.$
@@
For $\beta\in(0,2)$, $\operatorname{ES}_{\beta}$ is a strictly proper scoring rule, meaning that the expected energy score $\mathbb{E}_{\mathbf{X}}[\operatorname{ES}_{\beta}(\mathbb{P},\mathbf{X})]$ is maximized if and only if $\mathbf{X}\sim\mathbb{P}$, see §˜C.1 for details on probabilistic forecasting and scoring rules. Combined with its computational simplicity, this property makes the negative expected energy score a popular loss function for distributional regression (Bouchacourt et al., 2016; Shen & Meinshausen, 2024b, a; De Bortoli et al., 2025; Shen et al., 2025). It also directly implies the following corollary.
@@
The objective in (5.5) is minimized if and only if $\mathbb{P}_{h}=\widehat{\mathbb{P}}_{e\to h}$ for all $e,h\in[M]_{0}$.
Training. Since, in practice, we do not have access to the true distributions, we approximate expectations w.r.t. $\mathbb{P}_{e}$ and $\mathbb{P}_{h}$ in (5.6) with Monte Carlo estimates based on mini-batches of (2.1), and use a single draw of $\bm{\varepsilon}$} to approximate expectations w.r.t. $\mathbb{Q}_{\bm{\varepsilon}}$. We then optimise (5.5) w.r.t. $\widehat{\mathbf{W}}$ and the parameters of $\widehat{\mathbf{g}}$ and $\widehat{\mathbf{f}}$ using stochastic gradient descent (Robbins & Monro, 1951; Kingma & Ba, 2015).
@@
Prediction. To sample from the predicted distribution for a new perturbation label $\mathbf{a}_{\mathrm{test}}$, we use our model to compute the synthetic perturbed test latents $\widehat{\mathbf{z}}^{\mathrm{pert}}_{e\to\mathrm{test},i}$ for all $e\in[M]_{0}$ and all $i\in[N_{e}]$ via (5.2), and then sample synthetic test observations $\widehat{\mathbf{X}}_{e\to\mathrm{test},i}$ according to (5.3). Our estimate of $\mathbb{P}_{\mathrm{test}}$ is a convex combination of the domain-specific marginal perturbation distributions $\widehat{\mathbb{P}}_{e\to\mathrm{test}}$ induced by our model for $\mathbf{a}_{\mathrm{test}}$, i.e., the empirical version of
$\widehat{\mathbb{P}}_{\mathrm{test}}=\sum_{e\in[M]_{0}}\omega_{e}\widehat{\mathbb{P}}_{e\to\mathrm{test}}.$ (5.7)
with weights $\omega_{e}\geq 0$ such that $\sum_{e}\omega_{e}=1$ to account for prior knowledge or differing sample sizes across domains; common approaches include uniform weighting ($\omega_{e}\!=\!\frac{1}{M+1}$) or using only control observations ($w_{e}=0$ for $e\neq 0$).
Full algorithm and extensions. Due to space constraints, we present only the core PDAE architecture here; for additional loss components which may help in practice but are not used in our simulations (a reconstruction objective, latent prior, and sparsity regularization), a complete algorithm, and a discussion of goodness of fit testing, see Appx.˜B.
@@
Table 1: Results on synthetic 2D data. We report average performances over 7 ID and 14 OOD test cases; shown are the mean $\pm$ one standard deviation over 5 random seeds. Best results are highlighted in bold. Theoretical guarantees for PDAE rely on distributional overlap in the latent space (ID); no method is expected to perform well on OOD test cases.
@@
Pool All 0.855 $\pm$ .005 0.557 $\pm$ .004 0.441 $\pm$ .001 1.908 $\pm$ .004 1.266 $\pm$ .004 1.298 $\pm$ .002
Pseudobulking 0.596 $\pm$ .004 0.716 $\pm$ .006 0.382 $\pm$ .002 1.130 $\pm$ .005 0.965 $\pm$ .006 0.894 $\pm$ .004
Linear Regression 0.069 $\pm$ .001 0.181 $\pm$ .003 0.131 $\pm$ .003 0.536 $\pm$ .010 0.808 $\pm$ .012 0.524 $\pm$ .007
CPA 0.600 $\pm$ .239 1.152 $\pm$ .287 0.516 $\pm$ .165 1.538 $\pm$ .261 1.885 $\pm$ .241 1.132 $\pm$ .171
PDAE (Ours) 0.001 $\pm$ .000 0.003 $\pm$ .001 0.012 $\pm$ .003 0.018 $\pm$ .006 0.029 $\pm$ .007 0.090 $\pm$ .017
6 Related work
Biological perturbation models. Our model from §˜3 is closely related to prior work on modeling combinations of genetic or chemical perturbations. CPA (Lotfollahi et al., 2023) employs an adversarial autoencoder for estimation, see §˜C.4 for details. SVAE+ (Lopez et al., 2023) and SAMS-VAE (Bereket & Karaletsos, 2023) instead fit a generative model by maximising the evidence lower bound (ELBO; a lower bound to the marginal likelihood) within the variational autoencoder (VAE) framework (Kingma & Welling, 2014; Rezende et al., 2014). This requires evaluating the likelihood $p(\mathbf{x}|\mathbf{z})$, and thus generally comes with parametric assumptions. In contrast, PDAE is based on classical (non-variational) autoencoders and is trained using the energy score, which only requires being able to sample from the stochastic decoder. Both SVAE+ and SAMS-VAE additionally assume that the effects of perturbations on the latents are sparse (cf. Appx.˜B). SAMS-VAE and CPA adopt the same compositional perturbation model as (3.1), wherein the effect of perturbations can be decomposed into a sum of effects of elementary perturbations, captured by the linear form $\mathbf{W}\mathbf{a}$. In contrast, SVAE+ extends work on identifiable VAEs with sparsely changing mechanisms (Lachapelle et al., 2022, 2024) and models each perturbation condition separately. None of these prior perturbation models provide theoretical identifiability or extrapolation guarantees.
Testing for and modeling interactions. Xu et al. (2024) propose tests for the hypthesis of no (nonlinear) interactions in latent space, based on the assumption that perturbations affect disjoint latents and given suitable data from both individual and pairs of perturbations. Adduri et al. (2025) and He et al. (2025) propose modeling interactions via attention in transformer-based architectures (Vaswani et al., 2017).
Causal approaches. Zhang et al. (2024) and Liu et al. (2025) consider perturbation extrapolation from a causal representation learning perspective. They model elementary perturbations as shift interventions and use a VAE-based approach, which, similar to PDAE, is trained primarily to maximise the similarity between real and simulated perturbation distributions. Schneider et al. (2025) instead model interventions directly at the level of measurements.
Identifiable representation learning for extrapolation. Saengkyongam et al. (2024, §4) consider an identifiability problem similar to the one we study in §˜4 and leverage affine identifiability to derive out-of-support extrapolation guarantees. However, their model and proof strategy differ from ours, e.g., by requiring the training support of perturbations $\mathbf{a}$ to contain a non-empty open subset of $\mathbb{R}^{K}$.
7 Simulation study
In this section, we investigate the empirical performance of PDAE through a controlled simulation with synthetic data.
Methods. We compare our approach with CPA (Lotfollahi et al., 2023) and the following baselines:
  • Pool all: output the distribution and mean obtained by pooling all training observations;
  • Pseudobulking: like Pool All but pool only training data arising from individual perturbations involved in the combination to be predicted (e.g., pool data from $(1,0,0)^{\top}$ and $(0,1,0)^{\top}$ to predict $\mathbf{a}_{\mathrm{test}}=(1,1,0)^{\top}$);
  • Linear regression: linearly regress the domain-specific means of observations $\bm{\mu}^{\mathbf{x}}_{e}\in\mathbb{R}^{d_{X}}$ on $\mathbf{a}_{e}$ and use the resulting model to predict the test mean $\bm{\mu}_{\mathrm{test}}^{\mathbf{x}}$ from $\mathbf{a}_{\mathrm{test}}$.
The former two were also used by Lotfollahi et al. (2023); we propose the latter as an additional (stronger) baseline. Since CPA and linear regression only predict a mean, in these cases, we apply the test mean shift to all observations in the reference condition to simulate the test distribution and use this to compute distributional similarity metrics.
Metrics. To assess the distributional fit between predicted and true (empirical) test distributions, we use the energy distance (ED; Székely & Rizzo, 2013), i.e., twice the normalized negative expected (marginal) energy score, and the maximum mean discrepancy (MMD; Gretton et al., 2012) with Gaussian kernel and bandwidth chosen by the median heuristic, see §˜C.2 for more on measures of distributional similarity. Since some methods are primarily designed for mean prediction, we also report the L2 norm of the difference between the predicted and true mean, $\left\lVert\bm{\mu}^{\mathbf{x}}-\widehat{\bm{\mu}}^{\mathbf{x}}\right\rVert$.
@@ 7 Simulation study
Data. For ease of visualization, we first consider the case of ${d_{Z}}\!=\!{d_{X}}\!=\!2$-dimensional latents and observations. The base distribution $\mathbb{P}_{\mathbf{Z}}$ is a zero-mean, isotropic Gaussian with standard deviation $\sigma\!=\!0.25$. We consider K=3K\!=\!3 elementary perturbations with associated shift vectors $\mathbf{w}_{1}\!=\!(1,0)^{\top}$, $\mathbf{w}_{2}\!=\!(0,1)^{\top}$, and $\mathbf{w}_{3}\!=\!(1,1)^{\top}$ and create M+1=4M\!+\!1\!=4 training domains with perturbation labels $\mathbf{a}_{0}\!=\!(0,0,0)^{\top}$, $\mathbf{a}_{1}\!=\!(1,0,0)^{\top}$, $\mathbf{a}_{2}\!=\!(0,1,0)^{\top}$, and $\mathbf{a}_{3}\!=\!(0,0,1)^{\top}$. In this case, the sufficient diversity condition (iii) used to establish identifiability in Thm.˜4.1 is satisfied. For testing, we distinguish two scenarios: in-distribution (ID) and out-of-distribution (OOD) test cases, depending on whether the perturbed latents resulting from $\mathbf{a}_{\mathrm{test}}$ lie mostly within or outside the support of the empirical distributions of perturbed training latents, respectively. We consider 14 randomly generated ID and OOD test cases each, see Fig.˜9 and Appx.˜D for a visualization and further details. To generate observations, we use the complex exponential $\mathbf{x}\!=\!\mathbf{f}(\mathbf{z})\!=\!\mathrm{e}^{z_{1}}(\cos z_{2},\sin z_{2})$ as a deterministic nonlinear mixing function.
@@
Experimental details. We generate $N_{e}=2^{14}$ observations for each domain. Both CPA and PDAE use $4$-hidden-layer MLPs with $64$ hidden units as encoders and decoders, a $\widehat{d_{Z}}=2$-dimensional latent space, and are trained for $2000$ epochs using a batch size of $2^{12}$. For PDAE, we set the energy score exponent $\beta\!=\!1$ and use a learning rate of $0.005$. For CPA, we choose a $1$-hidden-layer MLP with $64$ hidden units for the adversarial classifier. For the remaining hyperparameters, we randomly sample $100$ joint configurations and choose the one with the smallest MMD across $7$ ID test cases, see Appx.˜D for details. We report results for average performance across the remaining $7$ ID and $14$ OOD test cases over $5$ repetitions of randomly generating data and initializing, training, and evaluating models.
@@
Figure 3: Qualitative results. Shown are kernel density estimates of the training data (black); an ID test case (blue) with $\mathbf{a}^{\textsc{id}}_{\mathrm{test}}\!=\!(1,1,0)^{\top}$, which was not seen during training but induces the same perturbed latent distribution as $\mathbf{a}_{3}$; and an OOD test case (orange) with $\mathbf{a}_{\mathrm{test}}^{\textsc{ood}}\!=\!(1,0,1)^{\top}$. PDAE recovers an affine transformation—since condition (ii) of Thm.˜4.1 is not enforced, cf. §˜A.1—of the true latents (top row) leading to accurate distributional predictions for the training and ID test domains (bottom row). In the OOD case, the test distribution in latent space is identified but supported in a space that was mostly not seen during training (orange, top right). As a result, the corresponding decoder output does not accurately match the true OOD distribution (orange, bottom row).
@@
Results. The results are shown in Tab.˜1. For the ID test setting, PDAE performs best, achieving a near-perfect distributional fit. CPA outperforms the pooling and pseudobulking baselines (consistent with results reported by Lotfollahi et al. (2023) on gene perturbation data), but does substantially worse than linear regression and PDAE at both mean and distributional prediction. For the OOD test setting, all methods expectedly perform worse, with PDAE yielding the least bad performance,  [^2] We believe that PDAE benefits from implicit regularization in the network and that worse counterexamples can be constructed. see Fig.˜3 and its caption for details.
Figure 4: Comparison for varying signal-to-noise ratio.
@@
Noisy observations. We additionally consider a modified version of the aforementioned experiment with ${d_{X}}=10$ (instead of ${d_{X}}=2$) observed dimensions by concatenating ${d_{\epsilon}}=8$ additional dimensions of mean-zero Gaussian noise and investigate the behavior for increasing noise variance (i.e., decreasing signal-to-noise ratio), see §˜D.2 for details. As shown in Fig.˜4, PDAE outperforms other methods for low to medium noise levels, whereas for high noise levels, all methods except CPA exhibit similar performance.
8 Discussion, limitations, and future work
Necessity of Gaussianity. Gaussianity of $\mathbb{P}_{\mathbf{Z}}$ is sufficient to prove identifiability, but may not be necessary. Indeed, PDAE does not explicitly enforce any particular latent distribution, but still yields good performance. To impose a specific prior such as that from assumption (ii) of Thm.˜4.1, one can penalize deviations of the encoded basal state distributions from the prior, e.g., using the energy score as discussed in Appx.˜B. Based on preliminary results with Laplacian basal states, we hypothesize that our results can be generalized to other distributions from the exponential family.